Auniform elastic string has length “a1” when the tension is T1 and length a2 when the tension is T2. show the natural length of the string is (a2T1-a1T2)/(T1-T2).

  1. Auniform elastic string has length “a1”  when the tension is T1 and length a2 when the tension is T2. show the natural length of the  string is  (a2T1-a1T2)/(T1-T2).

Grade:11

1 Answers

Anjan
8 years ago
Let the coefficient of string be k and natural length be a0
Now,
T= -k(a-a0) ------------ (1)
T2 = -k(a2-a0) ------------- (2)
 
Now, divide eqn. (1) by eqn. (2)
 
We get T1(a2 – a0) = T(a- a0)
On solving the eqn. we get the required soln.